# suppSupplement

### v1

@@ Abstract
Figure 3: Illustrative examples for automated discovery of scattering setups. Target scattering behaviours and the corresponding irreducible graphs identified by our automated search. (a) Isolator. Only signal from mode 1 to 2 is transmitted. (b) Directional coupler. Multiple input signals are summed up and transmitted to the output port. Graph (I-III) are the smallest irreducible graphs for a directional coupler with two inputs. Graph (I) has the smallest number of couplings, graph (II) and (III) have the smallest number of parametric drives. The other identified 97 further irreducible graphs (not shown) require either more resources or are gauge transformations of the shown graphs. Graph (I) can be generalized to more inputs. (c) Directional amplifier. Input from port 1 is amplified by a factor of $\sqrt{G}$, the other direction is blocked. The device has minimal back-action and the noise is quantum-limited. The graphs (I-IV) are the simplest solution without making use of intrinsic loss on the port modes. Allowing intrinsic losses, graph (V) is the simplest graph. Analytical terms discovered for the cooperativities ${\cal C}_{ij}$, the intrinsic loss rescaled by the outcoupling (red arrow), and the flux $\Phi_{132}$ are indicated. All shown graphs are stable. The unstable solutions are further discussed in the Supplemental Material.
In the following, we demonstrate the successful application of our method to a selected number of illustrative examples.
As a first proof of concept, we apply our scheme to implement an isolator, i.e. a two-port device with perfect transmission from the input to the output and zero reverse transmission, see Fig. 3(a). Our method shows that the simplest solution involves a single auxiliary mode and a single complex coupling. The continuous optimization routine provides also the cooperativities ${\cal C}_{12}={\cal C}_{23}={\cal C}_{31}=1$ and the synthetic flux $\Phi_{123}=\pi/2$, recovering the solution used in [32, 33, 25, 2, 34]. In the special limiting case of an overdamped auxiliary mode, $\kappa_{3}\gg\kappa_{1},\kappa_{2}$, this mode acts as a Markovian bath mediating a dissipative interaction [35]. By adiabatic elimination, one can derive a generalization of the dissipative isolator discussed in [35], see Supplemental Material. It is always possible to translate the solutions discovered by our algorithm into schemes in the spirit of Ref. [35], with an engineered reservoir replacing the auxiliary modes. A closed formula for the corresponding Lindblad Master equation is given in the Supplemental Material.
@@
The aim of artificial scientific discovery is to create generalizable insights. We now demonstrate that the solutions systematically discovered using our scheme can inspire the discovery of more complex devices, e.g. involving an arbitrarily large number of ports. As an example, we consider a fully directional coupler. This is a device that combines signals from multiple input ports and transmits them with equal transmission amplitude to an output port. Here, we aim for a fully directional coupler with zero reflection and reverse-transmission amplitudes, cf Fig. 3(b). We leave the transmission amplitude $t$ as a free parameter.
@@
First, we use our scheme to find possible solutions for $N_{\rm in}=2$. It turns out that at least two auxiliary modes are required to realize a two-port device. In total, we find 100 different irreducible graphs that differ in the number and complexity of the underlying couplings. We select the graphs with the minimal number of couplings, Graph (I) with six couplings, and the graphs with the smallest number of non-reciprocal couplings, Graph (II) and (III), requiring seven couplings but only one parametric drive. The remaining irreducible graphs are either gauge transformations of these graphs or require more resources in terms of couplings and parametric drives.
@@
This scenario is an example where the automated discovery process has produced solutions that lead to generalizable insights, an important goal in artificial scientific discovery. Indeed, inspecting Graph (I), we note that it can be viewed as the combination of two isolator graphs. This inspires a general ansatz for an $N_{\rm in}$-port coupler: a graph that comprises an isolator building block for each input port, see the right panel in Fig. 3(b). We then use our continuous optimization routine to verify the validity of the ansatz and calculate cooperativities, fluxes, and the transmission $t$ for the first few $N_{\rm in}$, up to $N_{\rm in}=5$ corresponding to $N=11$ nodes. We note that for such large graphs, an exhaustive search of the graph space would be daunting. From the special solutions, we infer a general solution valid for arbitrary $N_{\rm in}$: All cooperativities between an input-port mode and the corresponding auxiliary mode take the same value, ${\cal C}_{\rm in,aux}=1$, the cooperativities between the output-port mode and the other modes take the value ${\cal C}_{\rm out,in/aux}=1/N_{\rm in}$, and the synthetic field fluxes are the same for all isolator-like loops, $\Phi_{\mathrm{in},\mathrm{out},\mathrm{aux}}=\pi/2$. Finally the transmission turns out to be $t=1/\sqrt{N_{\mathrm{in}}}$, implying that no noise from the auxiliary modes is injected into the output port.
@@
In many interesting problems, one might be trying to solve a whole class of tasks where one or several parameters in the target scattering matrix are allowed to vary continuously. This type of challenge can be solved using our algorithm by first identifying a valid graph using a fixed value of the parameter, then using continuous optimization to create a dataset of solutions for varying parameter values (but fixed graph structure), and finally attempting to derive closed analytical formulas for the cooperativities and fluxes by symbolic regression.
@@
To illustrate this, consider several recent works [35, 34, 36, 2, 37, 38] that have proposed schemes to realize on-chip directional quantum-limited amplifiers using multi-mode circuits and to replace bulky state-of-the-art ferrite-based devices. The target device has arbitrary gain $G$ and is fully directional, i. e. has zero back-reflection, and zero reverse-transmission, cf Fig. 3(c). Moreover, both at the input and the output ports the added noise should reach the so-called quantum limit, i.e. the fundamental limit set by the laws of quantum mechanics [39]. The latter requirements can be translated into constraints on the scattering matrix $S$ that depend on the temperatures of the input and output ports. For both input and output ports at zero-temperature we recover the solution used in [25, 2, 35, 34]. In this solution, the added noise in the output port is injected into the multi-mode system from the same port. An interesting open problem in this context is to find the simplest design for a fully directional amplifier that is still quantum-limited in the presence of a hot output port [38]. Liu et. al. [38] approached this problem with a hand-crafted approach, discovering two small architectures comprising two auxiliary modes and six couplings. However, applying our method we discovered several smaller architectures, see Fig. 3(c). For more details of our derivations and results, please see the Supplemental Material. The smallest solution comprises a single auxiliary mode, three couplings, and has an under-coupled output port, i. e. most of the signal is dissipated into a (zero-temperature) loss channel. While this setup is not power-efficient, it fulfills all the imposed goals and represents an instructive counter-intuitive solution, since one could naively expect that the loss of a large part of the signal would be incompatible with quantum-limited amplification.
Figure 4: Optomechanical circulator. (a) Target operation: Realize a circulator between optical modes ($\hat{a}$) by mediating the interaction using mechanical modes ($\hat{b}$). (b) No couplings allowed between optical modes: Architecture proposed in [40] (left) and irreducible graphs identified by our scheme (right). Two mechanical modes ($\hat{b}$) mediate the interaction between the optical modes ($\hat{a}$). (c) Irreducible graphs identified by our optimisation scheme if direct passive couplings between the optical modes are allowed. (d) Deviation between the scattering matrix and the target behavior for the left graph in (c). The deviation approaches 00 in the asymptotic limit of $C_{ab},C_{bb}\to\infty$ with $C_{ab}^{2}/C_{bb}=1$ (dashed white line).
@@
In many physical systems, an ideal target behavior can be reached only in the special limiting case of one or more system parameters that tend to infinity or zero, e.g. a strong coupling limit. Even though this ideal limit is strictly unattainable in practice, adopting it promotes simplification and conceptual understanding. How can we introduce such fruitful asymptotic analysis in the context of artificial scientific discovery? One key complication is that often multiple dimensionless coupling parameters $\mathcal{C}_{j}$ need to become infinite simultaneously. This is resolved by rephrasing the description in terms of ratios like $\mathcal{C}_{j}/\mathcal{C}_{k}$ (or even, more generally, $\mathcal{C}_{j}^{\alpha}/\mathcal{C}_{k}$, with a suitable exponent, depending on the situation). This ratio stays finite while $\mathcal{C}_{j},\mathcal{C}_{k}\rightarrow\infty$, and its numerical value can be found by the discovery algorithm. At the same time, the limit itself can be represented as an idealized building block.
Coming back to the discovery of scattering setups, a generic example of such a building block is a far-detuned bus mode which mediates purely coherent interactions between other modes. Both the bus-mode detuning $\Delta^{(b)}$ and the couplings to the neighboring modes $a$ tend to infinity. However, the ratios $\mathcal{C}^{2}_{ab}/\mathcal{C}_{bb}=|g_{ab}|^{4}/(\kappa_{a}\Delta^{(b)})^{2}$ stay finite, and it is these ratios that determine the scattering matrix after integrating out the bus mode. We can thus incorporate the far-detuned mode by adding these ratios as optimization parameters and incorporating the interaction they mediate into the Hamiltonian for the remaining modes (see Supplemental Material). When eventually translating back to a physical setup, we can freely choose the absolute value of the bus-mode detuning with the only constraint that it is much larger than the corresponding decay rate $\kappa_{b}$. There are plenty of physical examples where such a reasoning is useful, see e.g. [6, 5, 40, 36].

### v2

@@ II Optimisation scheme
Figure 4: Illustrative examples for automated discovery of scattering setups. Target scattering behaviors and the corresponding irreducible graphs identified by AutoScatter. (a) Isolator. Only signal from mode 1 to 2 is transmitted. (b) Directional coupler. Multiple input signals are summed up and transmitted to the output port. Graph (I-III) are the smallest irreducible graphs for a directional coupler with two inputs. Graph (I) has the smallest number of couplings, graph (II) and (III) have the smallest number of parametric drives. The other identified 97 further irreducible graphs (not shown) require either more resources or are gauge transformations of the shown graphs. Graph (I) can be generalized to more inputs (right panel). To realize a coupler with $N_{\mathrm{in}}$ inputs, one has to repeat the graph structure (bottom right, blue background) $N_{\mathrm{in}}$ times. Analytical terms discovered for the cooperativities and fluxes are indicated, the transmission turns out to be $t=1/\sqrt{N_{\mathrm{in}}}$. (c) Directional amplifier. Input from port 1 is amplified by a factor of $\sqrt{G}$, the other direction is blocked. The device has minimal back-action and the noise is quantum-limited. The graphs (I-IV) are the simplest solution without making use of intrinsic loss on the port modes. Allowing intrinsic losses, graph (V) is the simplest graph. Analytical terms discovered for the cooperativities $C_{ij}$, the intrinsic loss rescaled by the outcoupling (red arrow), and the flux $\Phi_{132}$ are indicated. All shown graphs are stable. The unstable solutions are further discussed in the Supplemental Material.
III A simple example: Isolator
As a first proof of concept, we apply AutoScatter to implement an isolator, i.e., a two-port device with perfect transmission from the input to the output and zero reverse transmission, see Fig. 4(a). Our method shows that the simplest solution has one auxiliary mode and a single complex-valued beamsplitter coupling. The solution has the cooperativities $C_{12}=C_{23}=C_{31}=1$ and the synthetic flux $\Phi_{123}=\pi/2$. In this case, the unique irreducible solution had already been discovered with hand-crafted approaches as in [35, 34, 31, 2, 33].
Since our solutions are valid for arbitrary decay rates, they also cover the special limiting case in which the auxiliary modes are fastly decaying and act as Markovian baths, mediating dissipative interactions [36]. In Supplemental Material C, we provide a mapping from the graph’s weights to the Lindblad Master equation of a reservoir engineering scheme. Coming back to the specific isolator example, this mapping allows us to recover the dissipative isolator discussed in [36].
In addition to finding the simplest solution, any valid reducible graph provided by our method can be used as an ansatz by an experienced researcher to find the analytical expression for the corresponding manifold of solutions. The numerical data can provide further insights to refine the initial ansatz, e.g. suggesting constraints in the graph parameters. As an example of this approach, we find the analytical expressions that describe (up to gauge transformations) all three-modes solutions of the isolator problem and generalize the solution presented in [35, 34, 31, 2, 33, 36]. By using the empirically observed constraints that all dimensionless detunings and all cooperativities assume the same values, we find a single analytical expression parametrized by the flux $\Phi_{123}$ ,
$C_{ij}=1+4\left(\frac{\Delta_{i}}{\kappa_{i}}\right)^{2},\quad\frac{\Delta_{i}}{\kappa_{i}}=\frac{1}{2}\tan(\Phi_{123}-\pi/2),$ (4)
for $j\neq i$, and $i,j=1,2,3$. We note for later discussion that after replacing the auxiliary mode with a third port mode, the same solution above describes the most general three-mode circulator.
IV Discovering generalizable design patterns
An important goal of scientific discovery is to create insights that can be used to derive more general solutions. A common example of such an insight is a design pattern discovered in solving a simple problem, but which can also be reused as a building block to solve more complex problems. Graphical representations naturally provide insights that can be exploited to create insights. For example, chemical graphs are used by the Dendral to interpret mass spectra [11]. As well, the authors of [20, 45, 46, 47] discovered a new graph-based scheme to create many-body entangled quantum states, which allows them to automatically discover graph solutions with a small number of particles and to generalize them to more particles.
In our setting, we demonstrate how our graph-based solutions can be generalized to realize devices with an arbitrary number of ports. This is an example of how the fact that the weights of AutoScatter solutions are typically unique or in a low-dimensional manifold, contributes to their interpretability. This appealing feature is achieved by focusing on irreducible graphs and by our rescaling in Eq. 2. This is also the reason why we have been able to derive analytical expressions from most of the numerical solutions provided by AutoScatter, see the following example and Sections V and VI.
Example: Fully directional coupler
@@ IV Discovering generalizable design patterns
As an example, we consider a fully directional coupler. This device combines signals from multiple input ports and transmits them with equal transmission amplitude to an output port. Here, we aim for a fully directional coupler with zero reflection and reverse-transmission amplitudes, see Fig. 4(b). We leave the transmission amplitude $t$ as a free parameter. We aim to find a general solution valid for an arbitrary number of input ports $N_{\rm in}$.
@@
First, we use AutoScatter to find possible solutions for $N_{\rm in}=2$. It turns out that at least two auxiliary modes are required to realize a two-port device. In total, we find 100 different irreducible graphs that differ in the number and complexity of the underlying couplings. We select the graphs with the minimal number of couplings, Graph (I), and the graphs that require the smallest number of parametric pumps, Graph (II) and (III). For an explanation of how to count the number of pumps, see Section VIII. The remaining irreducible graphs are either gauge transformations of these graphs or require more resources in terms of couplings and parametric pumps.
@@
Inspecting Graph (I), we note that it can be viewed as the combination of two isolator graphs. This inspires a general ansatz for an $N_{\rm in}$-port coupler: a graph that comprises an isolator building block for each input port, see the right panel in Fig. 4(b). We then use our continuous optimization routine to verify the validity of the ansatz and calculate the free parameters for the first few $N_{\rm in}$, up to $N_{\rm in}=5$ corresponding to $N=11$ nodes. For such large graphs, an exhaustive search would be daunting. From the solutions, we infer the weights for arbitrary $N_{\rm in}$: All cooperativities corresponding to the interaction between an input mode and its respective auxiliary mode take the same value, $C_{\rm in,aux}=1$, the cooperativities corresponding to the coupling between the output mode and other modes take the value $C_{\rm out,in/aux}=1/N_{\rm in}$. The synthetic field fluxes are the same for all isolator-like loops, $\Phi_{\mathrm{in},\mathrm{out},\mathrm{aux}}=\pi/2$. Finally, the transmission turns out to be $t=1/\sqrt{N_{\mathrm{in}}}$, implying that no noise from the auxiliary modes is injected into the output port.
V Solving a continuous class of problems
Many interesting physical problems are parameterized by continuous parameters, e.g., the temperature or the value of an external field. Human researchers typically attempt to tackle these problems by deriving analytical solutions that are naturally interpretable. This becomes even more interesting when applied to optimization problems where the objective is continuous, e.g., when optimizing a variational quantum algorithm [48] to realize a continuous set of quantum gates.
@@ V Solving a continuous class of problems
In our multi-mode circuit setting, one is often interested in a class of problems parameterized by one or more continuous parameters in the target scattering matrix. This type of challenge can be solved using our algorithm by first identifying a valid graph using a fixed set of values of the parameters. Then, we use the continuous optimization to create a dataset of solutions for varying parameter values (but fixed graph structure). Finally, we attempt to derive closed analytical formulas for the cooperativities and fluxes by symbolic regression. This scheme can also be used in similar optimization problems with continuous objectives as discussed above.
Example: Directional quantum-limited amplifier
@@
To illustrate this, consider several recent works [36, 31, 33, 37, 2, 38, 39] that have proposed schemes to realize on-chip directional quantum-limited amplifiers using multi-mode circuits to replace bulky state-of-the-art ferrite-based devices. The target device has arbitrary gain $G$ and is fully directional, i. e., has zero back-reflection, and zero reverse-transmission. Moreover, both at the input and the output ports the added noise should reach the so-called quantum limit, i.e., the fundamental limit set by the laws of quantum mechanics [40]. These requirements can be translated into constraints $f_{j}$ (see Eq. 3) on the scattering matrix $S$ that depend on the temperatures of the input and output ports (see Appendix H for more details). For both input and output ports at zero-temperature AutoScatter recovers the solution used in [31, 2, 36, 33] (not shown). In this solution, the added noise in the output port is injected into the multi-mode system from the same port. For this reason, the device will not be quantum limited if the output port is hot.
An interesting open problem in this context is to find the simplest design for a fully directional amplifier that is still quantum-limited in the presence of a hot output port [39], see Fig. 4(c). Liu et. al. [39] addressed this problem with a handcrafted approach, discovering two small architectures comprising two auxiliary modes and six couplings. However, applying AutoScatter we discovered several architectures, see graphs (I-IV) in Fig. 4(c), which require less couplings and can also be implemented with less parametric pumps (see Section VIII).
In our description in Section I we assumed for simplicity that the port modes’ only loss channel is the out-coupling to some waveguide. In general, port modes can be subjected to further losses, e.g., an intrinsic loss channel, which is not monitored. Using the loss rates of these additional loss channels as degrees of freedom in the continuous optimisation (see Appendix F for more details), AutoScatter finds graph (V) in Fig. 4(c). This solution comprises a single auxiliary mode, three couplings, and has an under-coupled output port, i.e., most of the signal is dissipated into a (zero-temperature) loss channel. While this setup is not power-efficient, it fulfills all the imposed goals and represents an instructive counter-intuitive solution, since one could naively expect that the loss of a large part of the signal would be incompatible with quantum-limited amplification.
To discover the graphs in Fig. 4(c), we run the optimization with a constant value for the gain factor $G$. In a second step, we use our continuous optimization to create a dataset of solutions for different values of $G$, while keeping the graph structure fixed. Using simple symbolic regression, we were able to derive closed analytical formulas for the cooperativities, and fluxes as well as the full scattering matrix for all discovered irreducible graphs, see Supplemental Material D and Fig. 4(c).
VI Discovering asymptotic solutions
@@
In many physical systems, an ideal target behavior can be reached only in the special limiting case of one or more system parameters that tend to infinity or zero, e.g., a strong coupling limit. Even though this ideal limit is strictly unattainable in practice, adopting it promotes simplification and conceptual understanding. How can we introduce such fruitful asymptotic analysis in the context of artificial scientific discovery? One key complication is that often multiple parameters have to assume ideal limits simultaneously and that the way these limits are approached determines the desired target behavior.
We first discuss a simple scenario featuring only two parameters. Consider the task of applying an ideal gate to a dissipative qubit by driving it with a pulse of appropriate amplitude $\lambda$ and duration $T$. During a gradient-based optimization run, the gate time $T$ approaches the asymptotic limit $T\to 0$ to reduce dissipative effects. At the same time, the amplitude $\lambda$ goes to infinity to compensate for the shorter pulse. Finally, the asymptotic behavior, e.g., $\lambda\approx\pi/T$ for a $\pi$ pulse, can be extrapolated from the data, see Fig. 5(a). In this approach, one can deduce both the scaling law $\lim_{T\to 0}\lambda T=\mathrm{const.}$ and the numerical value of the constant from the data. In quantum circuits consisting of multiple gates, one now has multiple amplitudes and gate times. Instead of optimizing those gate parameters directly, one can reformulate the optimization task assuming instantaneous pulses parametrized by the pulse area $\lambda T$. In this way, one builds the prior knowledge of the scaling law into the model and obtains the respective values for $\lambda T=\mathrm{const.}$ directly without requiring extrapolation. This instantaneous pulse concept can be reused as a building block and is naturally used to solve efficiently complex optimization tasks involving more than one qubit and many pulses of different types in a variational quantum circuit [48].
Figure 5: Asymptotic solutions in automated discovery. (a) Possible asymptotic behavior in a two-dimensional parameter space. The ideal behavior is achieved when $C_{1}$ tends to zero and $C_{2}$ to infinity while reaching asymptotically a curve. The color represents the loss on this curve, as indicated by the color bar. The gray arrows indicate possible paths in parameter space that a gradient-based optimization could take. In the driven lossy qubit example, the parameter $C_{1}$ corresponds to the gate time $T$, and $C_{2}$ to the drive amplitude $\lambda$. (b) Asymptotic behavior in higher-dimensional parameter spaces. In this case, the three parameters $C_{1}$, $C_{2}$, and $C_{3}$ should diverge while approaching a specific plane to achieve the ideal behavior. The loss on this plane is shown together with two, out of infinitely many, possible paths of approach to the ideal behavior.
Now let us consider a more general setting with three or more dimensionless parameters that reach an unattainable limit, see Fig. 5(b). As for the qubit problem, we would like to reparametrize the problem in a way that allows us to take the asymptotic limit and subsequently perform our optimization without requiring extrapolation. For the reparametrization, we need the scaling laws of all parameters in terms of one of the parameters reaching the unattainable limit. However, these scaling laws are a priori unknown. Moreover, in a less constrained problem, the parameters could approach the unattainable limit on trajectories with different asymptotic limits making it difficult to extrapolate the potentially infinite number of asymptotic solutions or even the scaling laws of the parameters, see Fig. 5(b).
To address these general challenges, we propose an approach based on asymptotic building blocks. This concept can be applied to a broad range of systems that can be represented as weighted graphs. Just like the instantaneous pulse building block for the previously mentioned example of variational circuits, our building blocks build into the graph model an assumption regarding the scaling laws of the parameters. Here, a building block is a node or group of nodes that can be added to the graph and is defined by the scaling laws for the corresponding weights. This approach allows us to combine different building blocks and graph topologies to test a variety of asymptotic ansatzes automatically.
To be more specific, we come back to the discovery of scattering setups. In our experience, an asymptotic solution is often required in the presence of experimental constraints that prevent the port modes from interacting directly [6, 5, 41, 37]. This scenario is widespread for hybrid devices including transducers [6, 5], isolators [41], and amplifiers [37]. One or more auxiliary bus modes (for example, in optomechanics mechanical modes) are needed to mediate the interactions between port modes. To achieve the ideal behavior, the cooperativities $C_{a_{j}b_{k}}$ of the couplings between the bus modes $\hat{b}_{k}$ and the port modes $\hat{a}_{j}$ and possibly also the dimensionless detunings of the bus modes $\sqrt{C_{b_{k}b_{k}}}$ must diverge. On the other hand, the ratios of cooperativities $C_{a_{j}b_{k}}/C_{a_{l}b_{k}}$ or $C_{b_{k}b_{k}}/C^{\alpha}_{a_{l}b_{k}}$ (with suitable exponents $\alpha$, depending on the situation) tend to a constant.
Figure 6: Optomechanical circulator. (a) Target operation: Realize a circulator between optical modes ($\hat{a}$) by mediating the interaction using mechanical modes ($\hat{b}$). (b) Architecture proposed in [41] when no direct couplings are allowed between the optical modes. Instead, the interaction is mediated by far-detuned mechanical modes. (c) Irreducible graphs identified by AutoScatter when no direct couplings are allowed between the optical modes (top) or when the couplings between two optical modes must be passive (bottom). The shown graphs have the smallest possible number of couplings. There exist more irreducible graphs which require more couplings (not shown). (d) Effective interaction between the optical modes after integrating out the far-detuned modes in any of our discovered graphs. This graph resembles a generalized form of the well known circulator scheme discussed in [2, 33, 34, 35, 31]. Panels (e) and (f) refer to the bottom left graph in (c). (e) Manifold of asymptotic solutions. Every point on the curve $(C_{a_{1}b}/\sqrt{C_{bb}})(C_{a_{2}b}/\sqrt{C_{bb}})=1$ represents an asymptotic solution which approaches the ideal behavior for $C_{bb}\to\infty$. The inset shows the convergence speed $r$ as a function of the ratio $C_{a_{1}b}/C_{a_{2}b}$ of the asymptotic parameters $C_{a_{1}b}/\sqrt{C_{bb}}$, and $C_{a_{2}b}/\sqrt{C_{bb}}$. (f) Cross-section of the square deviation loss ${\cal L}$ between the scattering matrix $S$ and the target matrix $S_{\rm target}$. We fix $C_{ab}\equiv C_{a_{1}b}=C_{a_{2}b}$. ${\cal L}\to 0$ in the asymptotic limit $C_{ab},C_{bb}\to\infty$ with $C_{ab}^{2}/C_{bb}=1$ (dashed white line).
We can gain intuition on these asymptotic solutions and incorporate the respective scaling law as a building block in AutoScatter by integrating out one of the modes ($\hat{b}$ or $\hat{a}$ depending on the situation). In this way, we map the original graph to a smaller effective graph whose parameters depend only on the appropriate ratios of cooperativities in the asymptotic limit. This allows us to take explicitly the asymptotic limit and later obtain the asymptotic ratios using the continuous optimization of AutoScatter. For example, consider a set of far-detuned auxiliary bus modes $b_{k}$ as in [41]. The bus modes are only virtually occupied and, thus, mediate a purely coherent interaction between two or more port modes $a_{j}$. In this case, we can integrate them out to obtain an effective graph for the port modes $a_{j}$. In the new graph, the effective cooperatives $C^{\rm eff}_{jl}=C_{a_{j}b_{k}}C_{b_{k}a_{l}}/C_{b_{k}b_{k}}$ depend only on the asymptotic parameters $C_{a_{j}b_{k}}/\sqrt{C_{b_{k}b_{k}}}$, see Supplemental Material E.
Currently, our approach is limited to using certain asymptotic building blocks, which have to be predefined by a human expert. The same expert-aided approach could be adopted in other artificial scientific discovery problems, e.g., featuring a varying number of masses or charges of very different sizes. An important goal for future research is to automatically identify such limits and automatically define such new concepts.
Example: Optomechanical circulator